Category

Prior year

Current year

Accounts payable

41,400

45,000

Accounts receivable

115,200

122,400

Accruals

16,200

13,500

Additional paid in capital

200,000

216,660

Cash

???

???

Common Stock @ par value

37,600

42,000

COGS

131,400

171,328.00

Depreciation expense

21,600

23,858.00

Interest expense

16,200

16,911.00

Inventories

111,600

115,200

Long-term debt

135,000

137,826.00

Net fixed assets

375,421.00

399,600

Notes payable

59,400

64,800

Operating expenses (excl. depr.)

50,400

60,690.00

Retained earnings

122,400

136,800

Sales

255,600

337,713.00

Taxes

9,900

18,866.00

1. What is the firm's net income for the current year? (explain how you found net income please)

2. What is the firm's current year net profit margin?

3. What is the firm's gross profit margin for the current year?

4. What is the firm's current year operating profit margin?

5. What is the entry for the current year's cost of goods sold on a common-sized income statement?

Please explain each question on what is being used and how you found them because I am really lost, thank you !
Answer #1

For the Question No. 1

TRADING AND PROFIT & LOSS A/C AS OF CURRENT YEAR
PARTICULARS AMOUNT PARTICULARS AMOUNT
TO INVENTORIES        1,11,600.00 BY SALES 3,37,713.00
TO COGS        1,71,328.00 BY INVENTORIES 1,15,200.00
TO GROSS PROFIT        1,69,985.00
       4,52,913.00 4,52,913.00
TO OPERATING EXPENSES           60,690.00 BY GROSS PROFIT 1,69,985.00
TO DEPRECIATION EXPENSES           23,858.00
TO INTEREST EXPENSES           16,911.00
TO NET INCOME           68,526.00
TOTAL 33 1,69,985.00 TOTAL 1,69,985.00

As you can see that Net Income is being calculated first we added all the sales and the inventories of right side then we minus it with the COGS and inventories at the starting of the year and then we also less the Operating expenes and then the Depriciation Expenses and Interest Expenses and then we got our Net Income.

For the Question No. 2

Net Income and Net Profit are both same things.

As Net Profit Margin is being calculated as =Net Profit/Sales*100

= 68526/337713*100=20.29%

Hence, Net Profit Margin is 20.29%

For the Question No. 3

Gross Profit Margin is being calculated as =Sales-COGS/Sales

=337713-171328/337713*100

=166385/337713*100=49.26%

Hence Gross Profit Margin is 49.26%

For the Question No. 4

Operating Profit Margin is being Calculated= Sales-(COGS+All General Expenes and Administrative Expenses)/Sales*100

=337713-(171328+60690+23858+16911)/337713*100

=337713-272787/337713*100=19.22%

Hence Gross Profit Margin is 19.22%

For the Question No.5

It will be shown as below in the Common Income Statement .

TRADING AND PROFIT & LOSS A/C AS Of CURRENT YEAR

PARTICULARS

AMOUNT

PARTICULARS

AMOUNT

TO INVENTORIES

       1,11,600.00

BY SALES

3,37,713.00

TO COGS

       1,71,328.00

BY INVENTORIES

1,15,200.00

And the Entry for the same will be

To Cost of Goods Sold………… $500

By Mr. XYZ(Account Payable)…….$ 500